OpenStudy (anonymous):

A museum worker randomly displays 8 historical writings in a row. What is the probability that she lines them up from the oldest to the most recent from left to right?

OpenStudy (anonymous):

not very likely is it?

OpenStudy (anonymous):

haha not really

OpenStudy (anonymous):

probability she picks the oldest first is \(\frac{1}{8}\) then the probability second oldest is chosen next is \(\frac{1}{7}\) (there are 7 left to choose from, one is oldest) and so on and so on

OpenStudy (anonymous):

so it is going to be \(\frac{1}{8}\times \frac{1}{7}\times \frac{1}{6}\times ...\) more succinctly written as \(\frac{1}{8!}\)